OpenStudy (anonymous):

Solve using the square root method. x2 - 16 = 0 A. {-8, 8} B. {-4, 4} C. {8} D. {4}

OpenStudy (anonymous):

x2 means 2*x?

myininaya (myininaya):

i think she meant x^2

myininaya (myininaya):

it does say use the square root property

OpenStudy (anonymous):

no she asked this just now

myininaya (myininaya):

oh i see

myininaya (myininaya):

this is a repeat isn't it meow?

OpenStudy (anonymous):

x^2 my baddd

OpenStudy (anonymous):

No its not

myininaya (myininaya):

you just ask this question

myininaya (myininaya):

refer to the thread before this one someone gave you answer but if you are looking for an explanation someone could provide that

myininaya (myininaya):

add 16 on both sides

myininaya (myininaya):

then take the square root of both sides

myininaya (myininaya):

don't forget to put plus or minus

myininaya (myininaya):

Try not to post the same questions back to back pretty please.

OpenStudy (anonymous):

So confused :(

OpenStudy (anonymous):

where

OpenStudy (anonymous):

the whole thing like the answer

jimthompson5910 (jim_thompson5910):

\[\Large x^2-16=0\] \[\Large x^2=16\] \[\Large x^2=\pm\sqrt{16}\] \[\Large x=\sqrt{16} \ \text{or} \ x=-\sqrt{16}\] \[\Large x=4 \ \text{or} \ x=-4\] So the solution set is {4, -4}